D Question 7 3 pts Firm EFG is growing fast, and will produce earnings of $0.81/share next year. These earnings will grow at 3% per year, probably forever. EFG also has a cost of capital of 11.4%. What is the firms P/E ratio? D Question 8 3 pts The P/E ratio on a stock market is 12. The underlying real earnings growth rate is 1.6%, at a constant long-run inflation rate of 2.7%, what is the firms nominal long-run cost of capital? Question 9 3 pts What is the net profit margin of this firm? Sales Dividends Market Value Earnings (NI) $131 $15 $291 $30Please answer them Carry out calculations to at least 4 decimal places. Enter percentages as whole numbers. Example: 3.03% should be entered as 3.03. Do not include commas or dollar signs in numerical answers.
Answer #1

Question 7

price of the stock=earnings next year/(cost of capital-growth rate)

=0.81/(11.4%-3%)

=9.6429

firm's P/E ratio=9.6429/0.81

=11.9048 or 11.90